# Covariate Adjusted Response Adaptive Design with Delayed Outcomes

### v1

@@ 4 Theoretical investigation
This section provides theoretical justifications for the proposed forward-looking CARA design. To begin, we present the semiparametric efficiency bound for estimating average treatment effects in the presence of response delay in multi-stage adaptive experiments. We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect.
Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the unknown delay mechanism in each stage using a conservative rule. Encouragingly, we are able to justify Condition 1 in this setting under mild primitive assumptions. Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle-optimal allocation, as this would require perfect knowledge of the delay mechanism. Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method. Under conservative extrapolation, we show that the optimized asymptotic variance is guaranteed to weakly decrease as the experiment progresses.
Our main results are developed under the general setting of multi-valued treatments, that is, there can be more than two treatment arms. Therefore, we will adopt a more general notation for the estimated treatment effects. As before, let $\hat{\mu}_{T}(x,a)$ denote the estimated mean potential outcome for treatment arm $a$ and subgroup $x$. Then the estimated subgroup treatment effect is $\hat{\tau}_{T}(x,a^{\prime},a)=\hat{\mu}_{T}(x,a^{\prime})-\hat{\mu}_{T}(x,a)$, and the estimated average treatment effect is $\hat{\tau}_{T}(a^{\prime},a)=\sum_{x\in\mathcal{X}}\hat{p}_{T}(x)\hat{\tau}_{T}(x,a^{\prime},a)$.
4.1 Notation, assumptions and semiparametric efficiency bound
We begin by recalling our adaptive experiment setting: the experiment consists of $T$ stages, and in each stage $t=1,2,\dots,T$, $n_{t}$ participants are enrolled. Let $N_{t}=\sum_{s=1}^{t}n_{s}$ represent the cumulative sample size at the end of stage $t$, with $N=N_{T}$ as the total sample size. We use the notation $S_{i}=t$ to indicate that individual $i$ is enrolled in stage $t$.
For each individual, we observe some baseline covariates, $X_{i}\in\mathcal{X}$. The observed outcome variable is denoted by $Y_{i}$, and it is related to the potential outcomes through the relation $Y_{i}=\sum_{a\in\mathcal{A}}\mathds{1}_{(A_{i}=a)}Y_{i}(a)$, where $A_{i}$ denotes the actual treatment assigned. Our first assumption imposes standard regularity conditions on the covariates and the potential outcomes.
@@ 4.1 Notation, assumptions and semiparametric efficiency bound
(i) The covariates and potential outcomes, $\{X_{i},\{Y_{i}(a):a\in\mathcal{A}\}\}$, are independently and identically distributed across $i=1,2,\dots,N$. (ii) The covariates have finite support, that is, $|\mathcal{X}|<\infty$. Let $p(x)=\mathbb{P}[X_{i}=x]$, then $\min_{x\in\mathcal{X}}p(x)>0$. (iii) The potential outcomes have finite fourth moments: $\max_{x\in\mathcal{X}}\max_{a\in\mathcal{A}}\mu_{4}(x,a)<\infty$.
The next assumption addresses the delay mechanism. Recall that the delay is represented by the random variable $D_{i}$, which indicates the number of stages after which the experimenter can observe $Y_{i}$. By our definition, $Y_{i}$ is observed at the end of stage $t$ if and only if $D_{i}+S_{i}\leq t$.
@@
Finally, as we consider multi-stage experiment settings, we are in an asymptotic regime in which $T$ is fixed, and $n_{t}\to\infty$.
@@
We are now ready to present the semiparametric efficiency bound for average treatment effect estimation in the oracle setting, where there is perfect knowledge of the delay mechanism, and treatments are randomly assigned conditional on the covariates. This efficiency bound serves as the foundation and starting point for our proposed adaptive experimental design, as our algorithm minimizes a feasible version of the efficiency bound, constructed using accrued information.
@@
Let Assumptions 13 hold. In addition, assume (i) $S_{i}$ are independently and identically distributed with $\mathbb{P}[S_{i}=t]=r_{t}$; (ii) the treatment assignment probabilities, $e_{t}(a|x)=\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t]$, are bounded away from 0 and 1. Then the efficient influence function for estimating $\mu_{1}(a)$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$ $=\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\mu_{1}(X_{i},a)-\mu_{1}(a).$
In addition, the efficiency influence function for estimating $\tau(a^{\prime},a)$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime},a)\ =\ \psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a^{\prime})-\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$
$=$ $\frac{\mathds{1}_{(A_{i}=a^{\prime},D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|X_{i})\rho(T-t|X_{i},a^{\prime})}\Big{(}Y_{i}(a^{\prime})-\mu_{1}(X_{i},a^{\prime})\Big{)}$
$\qquad\qquad-\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|X_{i})\rho(T-t|X_{i},a)}\Big{(}Y_{i}(a)-\mu_{1}(X_{i},a)\Big{)}+\tau(X_{i},a^{\prime},a)-\tau(a^{\prime},a),$
@@
$\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}$ $+\frac{\sigma^{2}(x,a)}{\sum\limits_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}+\big{(}\tau(x,a^{\prime},a)-\tau(a^{\prime},a)\big{)}^{2}\Bigg{]}.$
@@ 4.2 Asymptotic properties of treatment effects estimates
In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotically normal distribution. The main conclusions of this section rely on a high-level condition regarding the consistency of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Baldi Antognini and Zagoraiou, 2015; Hu et al., 2015). In line with existing literature, the consistency of optimized treatment allocation rules is often needed to establish statistical guarantees of various adaptive designs. We recall that the treatment assignment probabilities are denoted by $\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t,\mathcal{H}_{t-1}]=\hat{e}^{*}_{t}(a|x)\in\mathcal{H}_{t-1}$, which are allowed to depend on historical data.
We begin with a high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
@@
We then present a useful lemma, which characterizes an asymptotically linear representation of the estimated mean potential outcomes.
Lemma 1 (Asymptotically linear representation for estimated mean potential outcomes).
@@
$\sqrt{N}\big{(}\hat{\mu}_{1,T}(x,a)-\mu_{1}(x,a)\big{)}$
$\quad=\frac{1}{p(x)\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\mathds{1}_{(X_{i}=x,A_{i}=a,D_{i}+S_{i}\leq T)}\big{(}Y_{i}-\mu_{1}(x,a)\big{)}+o_{\mathrm{p}}(1).$
Theorem 2 (Asymptotic normality of estimated subgroup treatment effects).
Let Assumptions 13 hold. Then under Condition 1,
$\sqrt{N}\big{(}\hat{\tau}_{T}(x,a^{\prime},a)-\tau(x,a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}\Big{(}0,\frac{1}{p(x)}\Big{(}\frac{\sigma^{2}(x,a^{\prime})}{\sum_{t=1}^{T}r_{t}e_{t}(a^{\prime}|x)\rho(T-t|x,a^{\prime})}+\frac{\sigma^{2}(x,a)}{\sum_{t=1}^{T}r_{t}e_{t}(a|x)\rho(T-t|x,a)}\Big{)}\Big{)},$
and
$\sqrt{N}\big{(}\hat{\tau}_{T}(a^{\prime},a)-\tau(a^{\prime},a)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}(0,\mathsf{V}),$
where $\mathsf{V}$ is defined in Theorem 1.
Before discussing statistical inference, we compare this theorem to the semiparametric efficiency bound established in the previous subsection. Specifically, while the two results, Theorems 1 and 2, yield the same variance formula, they are conceptually very different. The semiparametric efficiency bound is derived under a fixed treatment assignment regime, which we then use to guide our experimental design algorithm. Specially, our algorithm aims to minimize this efficiency bound. However, from an ex ante perspective, it may be unclear why this is a useful exercise, as it is not immediately apparent how to achieve this efficiency bound. Encouragingly, Theorem 2 helps “close the loop” by showing that, under mild regularity conditions, the average treatment effect estimator achieves the efficiency bound, thereby justifying our design objective. To conclude this subsection, the result below establishes valid statistical inference, where the variance estimator, $\hat{\mathsf{V}}_{T}$, is defined at the end of Section 3.
@@ 4.3 Convergence of the optimized treatment allocation
Previously we have shown that the treatment effect estimators are consistent and asymptotically normally distributed. These results build on the high-level condition that the empirical treatment allocation, $\hat{e}_{t}^{*}(\cdot)$, will converge in large samples (Condition 1). In this subsection, we further investigate the theoretical properties of the proposed experimental design—specifically, solved from (6)—and show that it satisfies Condition 1, thereby tying all loose ends. To give a road map, we will first show that the asymptotic analogs of our design algorithm have unique solutions. In this process, we will carefully distinguish the oracle problem of minimizing the semiparametric efficiency bound (which requires perfect knowledge of the delay mechanism) from the problem of sequential optimization using only accrued knowledge (which requires extrapolating the delay mechanism). Next, we show that the empirically optimized treatment allocation converges to some large-sample limit. To conclude this subsection, we will present a bound on efficiency loss due to extrapolating the delay mechanism.
@@
In our theoretical investigation, we will focus on the conservative method for estimating the delay mechanism. This allows us to provide simple and easy-to-interpret conditions, hence avoiding lengthy discussions. Despite our focus on one particular approach in this subsection, we show in our simulations that all three methods (conservative, optimistic, and neutral) deliver valid inference for the desired causal effects and more effectively achieves the design objectives compared to classical approaches.
To start, we recall that $\mathsf{V}=\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ denotes the efficiency bound. Its minimizers will be denoted by
$\{e_{t}^{*}(\cdot),1\leq t\leq T\}=\operatorname*{argmin}_{e_{1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta.$
As the semiparametric efficiency bound is specific to two treatment values, $a$ and $a^{\prime}$, we impose the constraints that the treatment probabilities in these two arms are bounded from below. However, it is worth mentioning that one can easily incorporate additional constraints.
@@
As we have discussed, the oracle optimal solution $\{e_{t}^{*}(\cdot),1\leq t\leq T\}$ is not achievable in general, as it builds on perfect knowledge of the delay mechanism. In any realistic experimental setting, however, this information can only be learned sequentially. In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|\cdot)$ for $d<t$ using information at the end of stage $t$. For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism. Following our discussion in Section 2, we define
$\rho^{\dagger}_{t}(d|\cdot)$ $=\min\big{\{}\rho(d|\cdot),\rho(t-1|\cdot)\big{\}}\ \text{for $d=1,2,\dots$}.$
@@
Alternatively, we write $\rho^{\dagger}_{t}(d|\cdot)=\rho(d|\cdot)$ for $d<t$, and $\rho^{\dagger}_{t}(d|\cdot)=\rho(t-1|\cdot)$ for $d=t,t+1,\dots$. It is helpful to compare this definition with Assumption 2. There, we assumed that the delay mechanism is time-homogeneous, so that $\rho(\cdot)$ is not indexed by $t$. On the other hand, $\rho^{\dagger}_{t}(\cdot)$ arises due to our imperfect knowledge about the delay mechanism and the conservative extrapolation employed, and as a result it will depend on a specific stage. It is not the true delay mechanism, but it is the asymptotic analogue of $\hat{\rho}_{t}(\cdot)$, which we used in our optimization algorithm.
To study convergence property of $\hat{e}^{*}_{t}(\cdot)$, we will make a definition first. Let $e^{\dagger}_{1}(a|\cdot)=e^{\dagger}_{1}(a^{\prime}|\cdot)=1/2$. Then for $t=2,3,\dots,T$, define recursively that
$e^{\dagger}_{t}(\cdot)=\operatorname*{argmin}_{e_{t}(\cdot)}\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t}(\cdot),e_{t+1}(\cdot),\dots,e_{T}(\cdot)),\ \text{subject to }e_{t}(a|\cdot),e_{t}(a^{\prime}|\cdot)\geq\delta,$
@@
$\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,a^{\prime})}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}^{\dagger}(a^{\prime}|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a^{\prime})e_{\ell}(a^{\prime}|x)}$
$+\frac{\sigma^{2}(x,a)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}^{\dagger}(a|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,a)e_{\ell}(a|x)}\Bigg{]}.$
Our first main result is to show that both $e^{*}_{t}(\cdot)$ and $e^{\dagger}_{t}(\cdot)$ are unique. To this end, the following assumption.
@@
For any $d\neq d^{\prime}$, ${\rho(d|x,a^{\prime})}/{\rho(d|x,a)}\neq{\rho(d^{\prime}|x,a^{\prime})}/{\rho(d^{\prime}|x,a)}$.
We remark that this assumptions helps establish the uniqueness of $e^{*}_{t}(\cdot)$. However, it is still possible to have ${\rho_{t}^{\dagger}(d|x,a^{\prime})}/{\rho_{t}^{\dagger}(d|x,a)}={\rho_{t}^{\dagger}(d^{\prime}|x,a^{\prime})}/{\rho_{t}^{\dagger}(d^{\prime}|x,a)}$. Fortunately, this is easily resolved by a tie-breaking rule. See the Supplementary Materials for a specific recommendation and additional discussions.
@@
Let Assumptions 14 hold. Then both $e_{t}^{*}(\cdot)$ and $e_{t}^{\dagger}(\cdot)$ are unique. In addition, Condition 1 holds with $\hat{e}_{t}^{*}(\cdot)\overset{\mathrm{p}}{\to}e_{t}^{\dagger}(\cdot)$.
Before closing this section, we provide two insights regarding the variance minimization problem. The first result suggests that as more information about the delay mechanism becomes available in the adaptive experiment, the optimized asymptotic variance never increases. While this result may seem quite natural, we remark that it does rely on the specific extrapolation we employed for imputing the delay mechanism, and to be specific, the conservative method we discussed earlier.
@@
Let Assumptions 14 hold. Define $\mathsf{V}^{*}$ as the minimized oracle asymptotic variance. Then for all $t=1,2,\dots,T-1$,
$\min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq$ $\min_{e_{t+2}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t+1}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \geq\ \mathsf{V}^{*}.$
Denote by $\mathsf{V}$ the asymptotic variance of the estimated average treatment effect. Then $\mathsf{V}\leq\mathsf{V}^{*}+O(1/T)$.

### v2

@@ 4 Theoretical investigation
This section provides theoretical justifications for the proposed forward-looking CARA design. To begin, we present the semiparametric efficiency bound for treatment effect estimation in a multi-stage setting with delayed outcomes. We then develop a general result on the consistency and asymptotic normality of the estimated average treatment effect. Our theoretical investigation then proceeds by examining the specific setting considered in Section 3, where we extrapolate the estimated delay mechanism in each stage using a conservative rule. Finally, we observe that in an experiment with only finitely many stages, it is generally not possible to achieve the oracle allocation, as this would require perfect ex ante knowledge of the delay mechanism. Therefore, the last part of our theoretical investigation addresses the efficiency loss resulting from extrapolating the delay mechanism in our design method.
4.1 Notation, assumptions and the efficiency bound
We begin by recalling our adaptive experiment setting: the experiment consists of $T$ stages, and in each stage $t=1,2,\dots,T$, $n_{t}$ participants are enrolled. Let $N_{t}=\sum_{s=1}^{t}n_{s}$ represent the cumulative sample size at the end of stage $t$, with $N=N_{T}$ being the total sample size. We use the notation $S_{i}=t$ to indicate that individual $i$ is enrolled in stage $t$. For each individual, we observe some baseline covariates $X_{i}\in\mathcal{X}$. The observed outcome variable is denoted by $Y_{i}$, and it is related to the potential outcomes through $Y_{i}=A_{i}Y_{i}(1)+(1-A_{i})Y_{i}(0)$, where $A_{i}\in\mathcal{A}$ denotes the actual treatment assigned and $\mathcal{A}=\{0,1\}$. Moments of the potential outcomes are $\mu_{s}(x,a)=\mathbb{E}[Y_{i}(a)^{s}|X_{i}=x]$ and $\mu_{s}(a)=\mathbb{E}[Y_{i}(a)^{s}]$; we omit the subscript whenever $s=1$.
Our first assumption imposes standard regularity conditions on the covariates and the potential outcomes.
@@ 4.1 Notation, assumptions and the efficiency bound
(i) The covariates and potential outcomes, $(X_{i},Y_{i}(0),Y_{i}(1))$, are independent and identically distributed across $i=1,2,\dots,N$.
(ii) The covariates have finite support (i.e., $|\mathcal{X}|<\infty$). Let $p(x)=\mathbb{P}[X_{i}=x]$, then $\min_{x\in\mathcal{X}}p(x)>0$.
(iii) The potential outcomes have finite fourth moments: $\max_{x\in\mathcal{X}}\max_{a\in\mathcal{A}}\mu_{4}(x,a)<\infty$.
The next assumption addresses the delay mechanism, where we adopt the notation $D_{i}$ to indicate the number of stages after which the experimenter can observe $Y_{i}$. By our definition, $Y_{i}$ is observed at the end of stage $t$ if and only if $D_{i}+S_{i}\leq t$. Also recall that the delay distribution is represented by $\rho(d|x,a)=\mathbb{P}[D_{i}\leq d|X_{i}=x,A_{i}=a]$.
@@
(ii) $\min_{x\in\mathcal{X}}\min_{a\in\mathcal{A}}\rho(0|x,a)>0$.
Finally, we consider multi-stage experiment settings.
@@
We are now ready to present the semiparametric efficiency bound for average treatment effect estimation in the oracle setting, where one has perfect knowledge of the delay mechanism and treatments are randomly assigned conditional on the covariates. This efficiency bound serves as the foundation and starting point for our proposed adaptive experimental design, as our algorithm minimizes a feasible version of the efficiency bound constructed using accrued information.
@@
Let Assumptions 13 hold. In addition, assume (i) $S_{i}$ are independent and identically distributed with $\mathbb{P}[S_{i}=t]=r_{t}$; (ii) the treatment assignment probabilities, $e_{t}(a|x)=\mathbb{P}[A_{i}=a|X_{i}=x,S_{i}=t]$, are bounded away from 0 and 1. Then the efficient influence function for estimating $\mu(a)$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|a)$ $=\frac{\mathds{1}_{(A_{i}=a,D_{i}+S_{i}\leq T)}}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|X_{i},a)e_{t}(a|X_{i})}\Big{(}Y_{i}(a)-\mu(X_{i},a)\Big{)}+\mu(X_{i},a)-\mu(a),$
where $a\in\mathcal{A}=\{0,1\}$. In addition, the efficient influence function for estimating $\tau$ is
$\psi(X_{i},A_{i},D_{i},Y_{i},S_{i})\ =\ \psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|1)-\psi(X_{i},A_{i},D_{i},Y_{i},S_{i}|0),$
@@
$\mathsf{V}=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}$ $+\frac{\sigma^{2}(x,0)}{\sum\limits_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}+\big{(}\tau(x)-\tau\big{)}^{2}\Bigg{]}.$
@@ 4.2 Asymptotic properties of treatment effects estimates
In this subsection, we demonstrate that the treatment effect estimator is consistent and admits an asymptotic normal distribution. The main conclusions of this section rely on a high-level condition regarding the convergence of the optimized treatment allocation rule, making these conclusions applicable to a wide range of CARA and a broad class of frequentist adaptive experimental design settings (Hu and Zhang, 2004; Antognini and Zagoraiou, 2015; Hu, Zhu and Hu, 2015). We begin with this high-level condition, which will be verified in the next subsection for our proposed forward-looking CARA design.
@@
Theorem 2 (Asymptotic normality of estimated treatment effects).
@@
$\sqrt{N}\big{(}\hat{\tau}_{T}(x)-\tau(x)\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}\Big{(}0,\frac{1}{p(x)}\Big{(}\frac{\sigma^{2}(x,1)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,1)e_{t}(1|x)}+\frac{\sigma^{2}(x,0)}{\sum_{t=1}^{T}r_{t}\rho(T-t|x,0)e_{t}(0|x)}\Big{)}\Big{)},$
and $\sqrt{N}\big{(}\hat{\tau}_{T}-\tau\big{)}\overset{\mathrm{d}}{\to}\mathcal{N}(0,\mathsf{V})$, where $\mathsf{V}$ is defined in Theorem 1.
It is helpful to compare this theorem to the semiparametric efficiency bound established in the previous subsection. While the two results, Theorems 1 and 2, yield the same variance formula, they are conceptually very different. The semiparametric efficiency bound is derived under a fixed treatment assignment regime, which we then use to guide our design algorithms—either by minimizing the efficiency bound for power improvement or by maximizing failure reduction subject to a variance upper bound. However, from an ex ante perspective, it may not be immediately clear why this is a useful exercise, as it is not obvious how the bound could actually be achieved. Encouragingly, Theorem 2 helps close the loop by showing that the asymptotic variance of the estimated treatment effect matches the bound in Theorems 1 under mild regularity conditions, thereby justifying our design objective. That said, it is worth clarifying that, in general, one cannot achieve the optimized oracle efficiency bound, an issue we discuss further below.
To conclude this subsection, the following result establishes valid statistical inference, where the variance estimator $\hat{\mathsf{V}}_{T}$ is defined at the end of Section 3.
@@ 4.3 Convergence of the optimized treatment allocation
Previously we have shown that the treatment effect estimators are consistent and asymptotically normally distributed. These results build on the high-level condition that the empirical treatment allocation, $\hat{e}_{t}^{*}(\cdot)$, will converge in large samples (Condition 1). In this subsection, we further investigate the theoretical properties of the proposed experimental design, specifically, solved from (10), and show that it satisfies Condition 1, thereby tying all loose ends. To give a road map, we will first show that the asymptotic analogs of our design algorithm have unique solutions. In this process, we will carefully distinguish the oracle problem of minimizing the semiparametric efficiency bound (which requires perfect knowledge of the delay mechanism) from the problem of sequential optimization using only accrued knowledge (which requires extrapolating the estimated delay mechanism). Next, we show that the empirically optimized treatment allocation converges to some large-sample limit. To conclude this subsection, we will present a bound on efficiency loss due to extrapolating the delay mechanism.
@@
In our theoretical investigation, we will focus on the conservative method for estimating the delay mechanism. This allows us to provide simple and easy-to-interpret conditions, hence avoiding lengthy discussions. Another appealing property of the conservative approach is that it can be understood from a “minimax” perspective, where we sequentially minimize the asymptotic variance under the worst-case scenario of the delay mechanism. Despite our focus on one particular approach in this subsection, we show in our simulations that all three methods (conservative, optimistic, and neutral) deliver valid inference.
@@
To start, we recall that $\mathsf{V}(e_{1}(\cdot),\dots,e_{T}(\cdot))$ is the objective function in (3). Its minimizers will be denoted by $e_{t}^{*}(a|x)$ for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$. As we have discussed, the oracle optimal solution is not achievable in general, as it builds on perfect knowledge of the delay mechanism. In any realistic experimental setting, however, this information can only be learned sequentially. In fact, even we set $n_{t}=\infty$, part of the delay mechanism can still remain unknown for any $t$: to be even more precise, one typically only learn $\rho(d|x,a)$ for $d<t$ using information at the end of stage $t$. For this reason, we have introduced feasible and concrete ways to extrapolate the delay mechanism. Following our discussion in Section 2, we define
$\rho^{\dagger}_{t}(d|x,a)$ $=\min\big{\{}\rho(d|x,a),\rho(t-1|x,a)\big{\}}.$
@@
It is helpful to compare this definition with Assumption 2. There, we assumed that the delay mechanism is time-homogeneous, so that $\rho(d|x,a)$ is not indexed by $t$. On the other hand, $\rho^{\dagger}_{t}(d|x,a)$ arises due to our imperfect knowledge about the delay mechanism and the conservative extrapolation employed, and as a result it will depend on a specific stage. It is not the true delay mechanism, but it is the asymptotic analogue of $\hat{\rho}_{t}(d|x,a)$, which we used in our optimization algorithm. To study convergence property of $\hat{e}^{*}_{t}(a|x)$, we will make a definition first. Let $e^{\dagger}_{1}(0|x)=e^{\dagger}_{1}(1|x)=1/2$. Then for $t=1,2,\dots,T$, define recursively that
$e^{\dagger}_{t+1}(1|\cdot)=1-e^{\dagger}_{t+1}(0|\cdot)=\operatorname*{argmin}_{e_{t+1}(\cdot)}\min_{e_{t+2}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot)),$
@@
$\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))$ $=\sum_{x\in\mathcal{X}}p(x)\Bigg{[}\frac{\sigma^{2}(x,1)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,1)e_{\ell}^{\dagger}(1|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,1)e_{\ell}(x)}$
$+\frac{\sigma^{2}(x,0)}{\sum\limits_{\ell=1}^{t}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,0)e_{\ell}^{\dagger}(0|x)+\sum\limits_{\ell=t+1}^{T}r_{\ell}\rho^{\dagger}_{t}(T-\ell|x,0)(1-e_{\ell}(x))}\Bigg{]}.$
Our first main result is to show that both $e_{t}^{*}(a|x)$ and $e_{t}^{\dagger}(a|x)$ are unique. To this end, we make the following assumption.
@@
For any $d\neq d^{\prime}$, ${\rho(d|x,1)}/{\rho(d|x,0)}\neq{\rho(d^{\prime}|x,1)}/{\rho(d^{\prime}|x,0)}$.
We remark that this assumptions helps establish the uniqueness of $e_{t}^{*}(a|x)$. However, it is still possible to have ${\rho_{t}^{\dagger}(d|x,1)}/{\rho_{t}^{\dagger}(d|x,0)}={\rho_{t}^{\dagger}(d^{\prime}|x,1)}/{\rho_{t}^{\dagger}(d^{\prime}|x,0)}$. Fortunately, this is easily resolved by a tie-breaking rule. See the Supplementary Materials for a specific recommendation and additional discussions.
@@
Let Assumptions 14 hold. Then for $a\in\{0,1\}$, $x\in\mathcal{X}$, and $t=1,2,\dots,T$, both $e_{t}^{*}(a|x)$ and $e_{t}^{\dagger}(a|x)$ are unique. In addition, Condition 1 holds with $\hat{e}_{t}^{*}(a|x)\overset{\mathrm{p}}{\to}e_{t}^{\dagger}(a|x)$.
Before closing this section, we provide two insights regarding the variance minimization problem. The first result suggests that as more information about the delay mechanism becomes available in the adaptive experiment, the optimized asymptotic variance always (weakly) decreases. While this result may seem natural, we note that it relies on the use of conservative extrapolation. As discussed earlier, conservative extrapolation can be understood from a minimax perspective, where the asymptotic variance is sequentially minimized under the worst-case scenario of the delay mechanism.
@@
Let Assumptions 14 hold. Define $\mathsf{V}^{*}$ as the minimized oracle asymptotic variance, and $\mathsf{V}$ to be the asymptotic variance of the estimated average treatment effect. Then (i) for all $1\leq s<t\leq T-1$,
$\mathsf{V}^{*}\ \leq\ \mathsf{V}\ \leq\ \min_{e_{t+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{t}^{\dagger}(e_{t+1}(\cdot),\dots,e_{T}(\cdot))\ \leq\ \min_{e_{s+1}(\cdot),\dots,e_{T}(\cdot)}\mathsf{V}_{s}^{\dagger}(e_{s+1}(\cdot),\dots,e_{T}(\cdot));$
and (ii) for any $\epsilon>0$ and $d_{\epsilon}$ with $\min_{x\in\mathcal{X},a\in\mathcal{A}}\rho(d_{\epsilon}|x,a)\geq 1-\epsilon$, $\mathsf{V}\leq\mathsf{V}^{*}+C(d_{\epsilon}T^{-1}+\epsilon)$, where $C$ does not depend on $T$, $\epsilon$ or $d_{\epsilon}$.